# HAMLOCK: HArdware-Model LOgically Combined attacK

### v1

@@ 3.2.1 Single Neuron HAMLOCK
Our single neuron attack is inspired by the principle of activation separation of the Data-Free Backdoor Attack (DFBA) [4]. However, compared to DFBA, our approach is more flexible and architecturally distinct: First, our attack only requires an activation separation on a single neuron, unlike DFBA, which must construct a complete layer-by-layer path; Second, our framework is more general, supporting both randomly chosen and optimized trigger patterns, whereas DFBA is limited to optimized triggers; Third, our activation separation threshold can be set to a flexible non-zero value while DFBA requires a threshold at or near zero.
@@
Our single neuron attack focuses on the first layer because the model weights in the first layer directly interact with the trigger pattern regions (i.e., nonzero elements of $\bm{m}$), providing attackers more leverage for activation separation without additional complexity. Next, we denote the weight and bias of a randomly selected neuron (our trigger neuron) in the first layer as $\bm{w}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$, respectively. Given an input $\bm{x}$, the activation of the trigger neuron under a monotonically increasing non-linear activation function $\sigma(\cdot)$ (e.g., ReLU) is computed as $a(\bm{x})=\sigma({\bm{w}}^{\top}\bm{x}+b).$
@@
Our goal is to leverage the trigger neuron activation to distinguish between clean and backdoor samples. Since backdoor inputs can be visually or semantically diverse, a key insight is to establish a strong and direct relationship between the trigger pattern $\bm{\delta}$ and the neuron’s activation. This encourages the model to focus its attention specifically on the regions where the trigger is embedded. To achieve this, we first modify the neuron’s weight to obtain $\bar{\bm{w}}$, which is set to 0 when $\bm{m}_{i}$, the $i$-th pixel in mask $\bm{m}$, is 0 and untouched otherwise.
@@
This ensures the activation of the neuron is solely dependent on the trigger pattern $\bm{\delta}$, not influenced by any background content in the backdoor sample $\bm{x}^{\prime}$. Next, we isolate the neuron’s response to the trigger pattern from the one to clean inputs by maximizing the trigger response. By leveraging the monotonicity of the activation function $\sigma(\cdot)$, we instead maximize the pre-activation value $a(\bm{x})$ by solving
$\max\ \bar{\bm{w}}^{\top}\bm{\delta},$ (1)
@@
where both the $\bar{\bm{w}}$ and/or $\bm{\delta}$ can be treated as the optimization variable(s). The solution to this maximization problem is achieved (ideally) when $\mathrm{sign}(\bar{\bm{w}})=\mathrm{sign}(\bm{\delta})$.
@@
However, if we choose to optimize over $\bm{\delta}$, we must additionally enforce that each component of $\bm{\delta}$ lies within a box constraint: $\delta_{i}\in[\delta_{i}^{l},\delta_{i}^{u}]$ for domains where inputs values are bounded (e.g., normalized images are in range of [0,1]). Under these conditions, the optimal solution $\bm{\delta}^{*}$ is given by:
$\bm{\delta}^{*}_{i}=\begin{cases}\delta_{i}^{l}&\text{if }\bar{w}_{i}<0,\\ \delta_{i}^{u}&\text{if }\bar{w}_{i}\geq 0,\end{cases}\quad\forall i.$ (2)
The solution above is conceptually similar to the analytical formulation used in the DFBA attack for changes in the first layer [4]. However, our generic formulation in Eq. (1) also allows modifying the weights while fixing the trigger patterns. When we optimize the weight $\bar{\bm{w}}$, we scale its magnitude by a factor $s$ to construct the modified weight $\bm{w}^{*}$ as:
$\bm{w}^{*}_{i}=\begin{cases}-s\bar{\bm{w}}_{i}&\text{if }\bar{\bm{w}}_{i}\cdot\delta_{i}<0,\\ s\bar{\bm{w}}_{i}&\text{if }\bar{\bm{w}}_{i}\cdot\delta_{i}>0,\\ 0&\text{if }\delta_{i}=0,\end{cases}$ (3)
@@
Importantly, maximizing the activation on backdoor samples (i.e., over $\bm{\delta}$ or $\bar{\bm{w}}$) alone is not sufficient. We must also ensure that activations on clean inputs remain below a chosen threshold $\tau$, enabling a clear separation between the two activations. Let $a^{*}=\bar{\bm{w}}^{\top}\bm{\delta}+b$ denote the activation on the optimized backdoor sample, with $b$ being an adjustable bias term. To ensure separation, we enforce the constraint:
$\bm{w}^{\top}\bm{x}+b\leq\tau,\quad\forall\bm{x}\in\mathcal{D},$
where $\tau$ is a threshold chosen such that clean inputs fall below it (e.g., $\tau\approx 0$). By appropriately tuning the bias $b$ and scale $s$ (for weight optimization in (3)), our method creates a clear activation separation around a threshold $\tau$, such that $a(\bm{x}^{*})$ is above $\tau$ for backdoored inputs while $a(\bm{x})$ is below it for clean inputs. Our parameter tuning allows $\tau$ to be either zero or a flexible non-zero value (see Figure˜4 in Appendix), which also helps to trivially evade pruning defenses [36]. The potential false detection arises if a clean input $\bm{x}$ contains a region $\bm{m}\cdot\bm{x}$ that is very similar to the trigger pattern $\bm{\delta}$. However, by carefully choosing or designing $\bm{\delta}$ to be visually or semantically distinct from natural images, such false positives are unlikely in practice.
@@ 3.2.2 Multi-Neuron HAMLOCK
Our single-neuron attack is effective and stealthy, evading state-of-the-art defenses (see Section˜5.2) with minimal hardware overhead. However, one potential limitation with this attack is that the search space of the (single) trigger neuron is not huge, and hence a dedicated defender could, in theory, identify the trigger neuron by inspecting individual neuron activations in the first layer with sophisticated detection methods. While no such defense currently exists, we cannot fully eliminate this possibility. Therefore, we next introduce a generalized attack that selects some random layers from the DNNs and alters their few neurons to signal the trigger presence. This attack is extremely hard to detect, as the possible combinations of different neurons and layers are countless.
@@
For the generalized attack, we typically do not target the first layer and hence no longer have direct access to the trigger regions for effective optimization. Nevertheless, we still seek to maximize the separation between the activations of clean and backdoor samples. We describe our attack using a fully connected layer for simplicity, but the principle applies directly to convolutional layers, as their core operation is also a weighted sum. For a given trigger neuron $j$ in layer $l$, it is connected to $N$ neurons from the previous layer l1l-1, and for a connected neuron $i$ in layer l1l-1, we denote their associated weight to $j$ as $w_{ji}$ and the produced activation as $a_{i}(\cdot)$.
Our objective is to optimize individual weights $w_{ji}$ of a trigger neuron $j$ so that the minimal activation difference between any backdoor and clean samples is maximized, which can be intractable as we have to iterate over many different combinations of backdoor and clean samples. To make the problem more tractable, we alternatively maximize the difference between the average activations between clean samples $\bm{x}$ and the corresponding backdoor samples $\bm{x}^{\prime}$, and this set of test samples are provided by the victim for hardware performance optimization, as described in Section˜2.2.2. For the multi-neuron attack, a backdoor sample $\bm{x}^{\prime}$ is obtained by adding a fixed trigger pattern onto the clean sample.
For a small number of $M$ test samples $\{\bm{x}_{1},...,\bm{x}_{M}\}$ provided by the victim, we let their average output activation $a_{i}(\cdot)$ from neuron $i$ (in layer l1l-1) for clean and backdoor samples as:
$\bar{a}_{i}(\bm{x})=\frac{1}{M}\sum_{m=1}^{M}a_{i}(\bm{x}_{m}),\quad\bar{a}_{i}(\bm{x}^{\prime})=\frac{1}{M}\sum_{m=1}^{M}a_{i}(\bm{x}^{\prime}_{m}).$
Then we modify the weight $w_{j}$ by optimizing:
$\max_{w_{j}}~~\sigma\bigg(\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\bm{x}^{\prime})\bigg)-\sigma\bigg(\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\bm{x})\bigg).$
@@
Since $\sigma(\cdot)$ is monotonically increasing, we again instead choose to maximize the pre-activation value difference:
$\max_{w_{j}}~~\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\bm{x}^{\prime})-\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\bm{x})=\sum_{i=1}^{N}w_{ji}\big(\bar{a}_{i}(\bm{x}^{\prime})-\bar{a}_{i}(\bm{x})\big).$ (4)
@@
And the heuristic solution to the optimization problem is simple: we simply set the new weight $w_{ji}^{\prime}$ as $w_{ji}^{\prime}=s\cdot w_{ji}\cdot\text{sign}(\bar{a}_{i}(\bm{x}^{\prime})-\bar{a}_{i}(\bm{x}))$, where $s>0$ is a scaling factor. Empirically, we find that a set of 100 test samples ($i.e.,M=100$) is sufficient to learn a robust separation that generalizes to all unseen clean and their corresponding backdoor test samples.
@@
The weight optimization strategy in Eq. (4) also informs our selection of optimal trigger neurons. Our goal is to choose neurons that maximize the activation separation while preserving the model’s clean accuracy. To select $k$ trigger neurons in each randomly chosen layer, we first identify a subset of viable candidates. This is done via neuron ablation: we set the activation of each neuron to 0 individually and measure the accuracy reduction on $M$ clean test samples. A negligible drop (e.g., within 2%) indicates that subsequent layers are insensitive to changes in that neuron’s output, making it a safe candidate for modification. From this pool of safe candidates, we then rank them by their absolute mean activation difference between the $M$ clean and backdoor samples, selecting the top-$k$ neurons as the final triggers for that layer.
@@
For trigger neurons across multiple layers, we optimize each layer sequentially, starting from the earliest chosen layer. Empirically, we find that a single neuron may not always separate the backdoor and clean activations perfectly; some clean activations might still cross the defined threshold. However, when we consider multiple trigger neurons in conjunction, the chance of a clean activation simultaneously exceeding all thresholds becomes extremely rare. This makes separating the backdoor and clean activations easy, as we only consider samples that cross the defined thresholds of all trigger neurons, which can be implemented efficiently in hardware.
@@ 3.3 Hardware Trojan Details
For the single-neuron or multi-neuron attacks with non-zero detection thresholds, the HT observes the exponent fields of the single or multiple trigger neuron activations. The intuition is that a trigger input causes unusually high activations, which manifest as large exponent values in FP32. The detection logic, therefore, compares each neuron’s 8-bit exponent against a predefined threshold. The trigger condition is asserted only when all monitored neurons exceed their corresponding thresholds. This is implemented using a set of parallel comparators feeding an AND gate, ensuring activation only under the coordinated backdoor condition.
@@
The bias injection can be realized in two ways. A hardcoded design directly embeds $b^{\prime}$ as a fixed constant at synthesis time. This minimizes hardware overhead and maximizes stealth, but lacks adaptability. Alternatively, a reconfigurable design stores $b^{\prime}$ in a small register, allowing it to be updated at runtime. While slightly more complex, this approach enables dynamic payload adjustment. For our experiments in Section˜5, we adopt the hardcoded design, as it is sufficient to guarantee reliable misclassification.
@@
Implementation of class-specific misclassification is strightforward. The hardware payload Trojan can be augmented with additional logic that monitors the model’s predicted output (e.g., the argmax logit index), which is implemented via a simple comparator circuit that observes which output neuron has the maximum logit value and maps it to a predetermined target label. The final class output can then be overridden by injecting a bias into the logit of the designated wrong class, based on the current prediction. While this approach adds moderate logic overhead compared to the class-agnostic case, it remains practical and maintains the modular design.
RTL Realization. The Trojan was implemented in synthesizable Verilog. A simplified multi-neuron trigger variant (code snippet) is shown below, where multiple neuron outputs are compared against threshold values and, upon satisfaction, a noise bias is added to the payload neuron’s input:
@@
end
This modular design ensures that both the trigger and payload logic of HTs incur negligible hardware overhead while maintaining deterministic activation behaviour.
Hardware Implementation Details. The trigger and payload circuits of HAMLOCK are implemented in synthesizable Verilog using a standard-cell-based commercial ASIC design flow. The synthesis was performed with Synopsys Design Compiler targeting the 45nm gscl45nm technology library under timing constraint (i.e, no impact on performance). This methodology enables accurate gate-level estimation of both area and power overheads introduced by HTs.

### v2

@@ 3.2.1 Single Neuron HAMLOCK
We introduce our model-level single neuron attack, which is inspired by the Data-Free Backdoor Attack (DFBA) [4], and we describe our key distinctions to DFBA at the end of this section. The single neuron attack focuses on the first layer because the model weights in the first layer directly interact with the trigger pattern regions (i.e., nonzero elements of $\boldsymbol{m}$), providing attackers more leverage for activation separation without additional complexity. Next, denote the weight and bias of a randomly selected trigger neuron in the first layer as $\boldsymbol{w}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$, respectively. Given an input $\boldsymbol{x}$, the activation of the trigger neuron under a monotonically increasing non-linear activation function $\sigma(\cdot)$ (e.g., ReLU) is computed as $a=\sigma({\boldsymbol{w}}^{\top}\boldsymbol{x}+b).$
@@
Our goal is to leverage the activation of neuron $k$ to distinguish between clean and backdoor samples. Since backdoor inputs can be visually or semantically diverse, a key insight is to establish a strong and direct relationship between the trigger pattern $\boldsymbol{\delta}$ and the neuron’s activation. This encourages the model to focus its attention specifically on the regions where the trigger is embedded. To achieve this, we first modify the neuron’s weight to obtain $\bar{\boldsymbol{w}}$, which is set to 0 when $\boldsymbol{m}_{i}$, the $i$-th pixel in mask $\boldsymbol{m}$, is 0 and untouched otherwise.
@@
This ensures the activation of neuron $k$ is solely dependent on the trigger pattern $\boldsymbol{\delta}$, not influenced by any background content in the backdoor sample $\boldsymbol{x}^{\prime}$. Next, we isolate the neuron’s response to the trigger pattern from the one to clean inputs by maximizing the trigger response. By leveraging the monotonicity of the activation function $\sigma(\cdot)$ (e.g., ReLU), we instead maximize the pre-activation value $a$ by solving
$\max\ \bar{\boldsymbol{w}}^{\top}\boldsymbol{\delta},$ (1)
@@
where both the $\bar{\boldsymbol{w}}$ and/or $\boldsymbol{\delta}$ can be treated as the optimization variable(s). The solution to this maximization problem is achieved (ideally) when $\mathrm{sign}(\bar{\boldsymbol{w}})=\mathrm{sign}(\boldsymbol{\delta})$.
@@
However, if we choose to optimize over $\boldsymbol{\delta}$, we must additionally enforce that each component of $\boldsymbol{\delta}$ lies within a box constraint: $\delta_{i}\in[\delta_{i}^{l},\delta_{i}^{u}]$ for domains where inputs values are bounded (e.g., normalized images are in range of [0,1]). Under these conditions, the optimal solution $\boldsymbol{\delta}^{*}$ is given by:
$\boldsymbol{\delta}^{*}_{i}=\begin{cases}\delta_{i}^{l}&\text{if }\bar{w}_{i}<0,\\ \delta_{i}^{u}&\text{if }\bar{w}_{i}\geq 0,\end{cases}\quad\forall i.$ (2)
When we instead optimize the weight vector $\bar{\boldsymbol{w}}$, we scale its magnitude by a factor $s$ to construct the modified weight $\boldsymbol{w}^{*}$ as:
$\boldsymbol{w}^{*}_{i}=\begin{cases}-s\bar{\boldsymbol{w}}_{i}&\text{if }\bar{\boldsymbol{w}}_{i}\cdot\delta_{i}<0,\\ s\bar{\boldsymbol{w}}_{i}&\text{if }\bar{\boldsymbol{w}}_{i}\cdot\delta_{i}>0,\\ 0&\text{if }\delta_{i}=0,\end{cases}$ (3)
@@
Importantly, maximizing the activation on backdoor samples (i.e., over $\boldsymbol{\delta}$ or $\boldsymbol{x}$) alone is not sufficient. We must also ensure that activations on clean inputs remain below a chosen threshold $\tau$, enabling a clear separation between the two activations. Let $a^{*}=\bar{\boldsymbol{w}}^{\top}\boldsymbol{\delta}+b$ denote the activation on the optimized backdoor sample, with $b$ being an adjustable bias term. To ensure separation, we enforce the constraint:
$\boldsymbol{w}^{\top}\boldsymbol{x}+b\leq\tau,\quad\forall\boldsymbol{x}\in\mathcal{D},$
where $\tau$ is a threshold chosen such that clean inputs fall below it (e.g., $\tau\approx 0$). By appropriately tuning the bias $b$ and scale $s$, our method creates a clear activation separation around a threshold $\tau$, such that $a(\boldsymbol{x}^{*})$ is above $\tau$ for backdoored inputs while $a(\boldsymbol{x})$ is below it for clean inputs. Our parameter tuning allows $\tau$ to be either zero or a flexible non-zero value (see Figure˜4 in Appendix), which also helps to trivially evade pruning defenses [37]. The potential false detection arises if a clean input $\boldsymbol{x}$ contains a region $\boldsymbol{m}\cdot\boldsymbol{x}$ that is very similar to the trigger pattern $\boldsymbol{\delta}$. However, by carefully choosing or designing $\boldsymbol{\delta}$ to be visually or semantically distinct from natural images, such false positives are unlikely in practice. We use $s=1$ for all settings in our experiments.
Distinction from prior work. The solution in Eq. (2) is conceptually similar to the analytical formulation used in the DFBA attack for changes in the first layer [4]. However, our key distinctions from DFBA are: 1) our attack is based on general root formulation in Eq. (1), which allows for both the weight and trigger optimization, while DFBA is less general and only considers trigger optimization. Second, our attack only modifies a single neuron for minimal footprint, while DFBA modifies neurons in all layers, making them ineffective against state-of-the-art defenses (Section˜5.2). Third, our activation separation threshold $\tau$ can be set to a flexible non-zero value, while DFBA requires a threshold at or near zero.
@@ 3.2.2 Multi-Neuron HAMLOCK
The single neuron attack is effective and stealthy (evades the state-of-the-art defenses, as shown in Section˜5.2), while posing minimal overhead for the hardware. However, in practice, there might be dedicated defenders who specifically inspect the individual neuron activations in the first layer, which might detect our attack. Therefore, we next introduce a generalized attack that selects $r$ random layers from the DNNs and alters few neurons to signal the trigger presence. This attack is extremely hard to detect, as the possible combinations of different neurons and different random layers are countless.
@@
Moving from the first layer, we no longer have direct access to the trigger regions for effective optimization, but we still seek to maximize the separation between the activations of clean and backdoor samples. We describe our attack using a fully connected (FC) layer for simplicity, but the principle applies identically to convolutional layers as the core operation is a weighted sum. For a given trigger neuron $j$ in layer $l$, it is connected to $N$ neurons from the previous layer l1l-1, with associated weight $w_{ji}$ and denote their activation values as $a_{i}$. Our objective is to optimize individual weights $w_{ji}$ of a trigger neuron $j$ to maximize the activation separation, ensuring any backdoored activation is greater than the clean one:
$\max_{w_{j}}~~\sigma\bigg(\sum_{i=1}^{N}w_{ji}a_{i}(\boldsymbol{x}^{\prime})\bigg)-\sigma\bigg(\sum_{i=1}^{N}w_{ji}a_{i}(\boldsymbol{x})\bigg).$
@@
Since $\sigma(\cdot)$ is monotonically increasing, we can equivalently maximize the pre-activation value difference:
$\max_{w_{j}}~~\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\boldsymbol{x}^{\prime})-\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\boldsymbol{x})=\sum_{i=1}^{N}w_{ji}\big(\bar{a}_{i}(\boldsymbol{x}^{\prime})-\bar{a}_{i}(\boldsymbol{x})\big).$
Ideally, we aim to maximize the activation separation for any arbitrary test sample $\boldsymbol{x}$, which can then be formulated as a max-min optimization problem as follows:
$\max_{w_{j}}\min_{\boldsymbol{x},\boldsymbol{\delta}}~~\sum_{i=1}^{N}w_{ji}\big({a}_{i}(\boldsymbol{x}^{\prime})-{a}_{i}(\boldsymbol{x})\big),~~~\boldsymbol{x}^{\prime}=\boldsymbol{x}\cdot(1-\boldsymbol{m})+\boldsymbol{m}\cdot\boldsymbol{\delta}.$ (4)
However, exactly solving the optimization problem above can be intractable. Therefore, we make two approximations to efficiently solve the problem in practice. First, we use a fixed trigger pattern $\boldsymbol{\delta}$ to obtain $\boldsymbol{x}^{\prime}=\boldsymbol{x}\cdot(1-\boldsymbol{m})+\boldsymbol{m}\cdot\boldsymbol{\delta}$. Then we replace the inner minimization of the activation separation with an average over a small set of $M$ clean samples $\{\boldsymbol{x}_{1},...,\boldsymbol{x}_{M}\}$ readily available to the attacker (Section˜2.2.2). This helps to simplify the max-min optimization into a single level optimization. Specifically, we compute the average output activation $\bar{a}_{i}(\cdot)$ on $M$ samples from neuron $i$ for clean and backdoor samples as:
$\bar{a}_{i}(\boldsymbol{x})=\frac{1}{M}\sum_{m=1}^{M}a_{i}(\boldsymbol{x}_{m}),\quad\bar{a}_{i}(\boldsymbol{x}^{\prime})=\frac{1}{M}\sum_{m=1}^{M}a_{i}(\boldsymbol{x}^{\prime}_{m}).$
and the max-min optimization now becomes:
$\max_{w_{j}}~~\sigma\bigg(\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\boldsymbol{x}^{\prime})\bigg)-\sigma\bigg(\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\boldsymbol{x})\bigg)$ (5)
Of course, future work could design more effective attacks by jointly optimizing $\boldsymbol{\delta}$ and $\boldsymbol{x}$ in Eq. (4).
@@
Second, when solving the maximization in Eq. (5), we use a fast heuristic to set the new weight $w_{ji}^{\prime}$ as $w_{ji}^{\prime}=s\cdot w_{ji}\cdot\text{sign}(a_{i}(\boldsymbol{x}^{\prime})-a_{i}(\boldsymbol{x}))$, where $s>0$ is a uniform scaling factor. Empirically, we find that a set of just 100 clean images and their corresponding backdoored versions is sufficient to learn a robust separation that generalizes to all unseen samples.
@@
However, problem in Eq. (5) is non-convex and hence, the initial values of $w_{j}$ matter. To improve attack effectiveness, we therefore select more potent trigger neurons from the $r$ random layers that achieve better activation separation after optimization while still preserving the model’s clean accuracy. To select $k$ more potent trigger neurons in a randomly chosen layer, we first identify a subset of viable candidates that do not hurt clean accuracy after alteration. This is done via neuron ablation: we set the activation of each neuron to 0 individually and measure the accuracy reduction on the $M$ selected clean test samples. A negligible drop indicates that subsequent layers are insensitive to changes in that neuron’s output, making it a safe candidate for modification. For our experiments in Section˜5, we set the individual accuracy drop threshold to 0.05% so that we can ensure the overall accuracy drop after all weight modifications remains within 3%. From the safe candidate pool, we then rank them by their absolute mean activation difference between the $M$ clean and backdoor samples, selecting the top-$k$ neurons as final trigger neurons.
@@
When we optimize over the $r$ randomly chosen layers, we optimize each layer sequentially, starting from the earliest chosen layer. Empirically, we find that a single neuron (i.e., $r=1,k=1$) may not always separate the backdoor and clean activations perfectly; some clean activations might still cross the separation threshold. However, when we consider multiple trigger neurons distributed across layers in conjunction, the chance of a clean activation simultaneously exceeding all thresholds becomes extremely rare. This makes separating the backdoor and clean activations easy, as we only consider samples that cross the defined thresholds of all trigger neurons as backdoors, which can be implemented efficiently in hardware. The complete optimization routine is summarized in Algorithm 1 in Appendix D. For our experiments, we set $r=3$ and $k=1$ for each of the randomly selected layers.
Distinction from prior work. Our final multi-neuron attack on an individual neuron shares a conceptual similarity to the heuristic approach in the handcrafted backdoor [21], as the design choices are natural for effective activation separation, but with three key distinctions: First, different from the proposed heuristic, our methodology is derived from a rigorous mathematical formulation that captures the essence of the optimal multi-neuron attack, and allows room for optimizing Eq. (4) better. Second, our attack is designed for sparsity and independence, requiring modification of only a few randomly selected layers ($r=3$) that are largely independent from each other and monitoring only a few neurons per layer ($k=1$). This contrasts to the entire modification with more neurons as in prior work, where the layers are chained and modifications in earlier layers determine the strategies in the subsequent layers, making the attack brittle. Third, the detection logic uses exponent bit comparisons and their conjunctions as a generalized local logic primitive. The handcrafted attack must merge all computations into the final logit layer to force a misclassification, which increases its reliance on global outputs and being detected by existing defenses [4].
@@ 3.3 Hardware Trojan Details
For the single-neuron attack with non-zero detection threshold or multi-neuron attack, the HT observes the exponent fields of the single or multiple trigger neuron activations. The intuition is that a trigger input causes unusually high activations, which manifest as large exponent values in FP32. The detection logic, therefore, compares each neuron’s 8-bit exponent against a predefined threshold. The trigger condition is asserted only when all monitored neurons exceed their corresponding thresholds. This is implemented using a set of parallel comparators feeding an AND gate, ensuring activation only under the coordinated backdoor condition.
@@
The bias injection can be realized in two ways. A hardcoded design directly embeds $b^{\prime}$ as a fixed constant at synthesis time. This minimizes hardware overhead and maximizes stealth, but lacks adaptability. Alternatively, a reconfigurable design stores $b^{\prime}$ in a small register, allowing it to be updated at runtime. While slightly more complex, this approach enables dynamic payload adjustment. For our experiments (Section˜5), we adopt the hardcoded design, as it is sufficient to guarantee reliable misclassification.
@@
Implementation of class-specific misclassification is straightforward. The hardware payload Trojan can be augmented with additional logic that monitors the model’s predicted output (e.g., the argmax logit index), which is implemented via a simple comparator circuit that observes which output neuron has the maximum logit value and maps it to a predetermined target label. The final class output can then be overridden by injecting a bias into the logit of the designated wrong class, based on the current prediction. It adds slightly higher logic overhead compared to the class-agnostic case, but maintains the modular design of HAMLOCK.
Robustness to build pipeline variability. HAMLOCK is robust to build-pipeline variability for two common scenarios. First, in dedicated accelerators for a fixed model, the mapping between software neurons and hardware compute units is established at design time and remains static. In this case, the HTs are directly integrated into the hardware neurons or datapath elements that implement the trigger software neurons. Consequently, variations in compiler flags, memory layout, or build-time optimizations do not affect the attack, as the software–hardware alignment is inherently guaranteed by construction and no additional coordination is required.
Second, in general-purpose accelerators supporting multiple models and runtime reconfigurations, software neurons are decoupled from physical hardware neurons. Here, the HTs are pre-implanted in a small subset of physical neurons or processing elements, and the attack schedules the affected software neurons to these Trojan-infected hardware resources for computation at runtime. Our threat model assumes the adversary is in control over the hardware vendor’s hardware-software stack, and this steering can be achieved reliably in the compilation process with minimal effort and remains robust against changes in optimization flags or memory layout.
In the rare event of misalignment, the attacker can automatically re-synthesize the software model to re-establish the desired mapping. To test new datasets and models, the only manual effort is to provide the trigger neuron location in the model to the automated process above.
RTL realization. The Trojan was implemented as a synthesizable register-transfer level (RTL) Verilog module. A simplified multi-neuron trigger variant (code snippet) is shown below, where multiple neuron outputs are compared against threshold values and, upon satisfaction, a noise bias is added to the payload neuron’s input:
@@
This modular design ensures that both the trigger and payload logic of HTs incur negligible hardware overhead while maintaining deterministic activation behavior, particularly when compared to existing HTs that require heavy computations (see Section˜6 for details).
Hardware implementation details. The trigger and payload circuits of HAMLOCK are implemented in synthesizable Verilog using a standard-cell-based commercial ASIC design flow. The synthesis was performed with Synopsys Design Compiler targeting the 45nm gscl45nm technology library under timing constraint (i.e., no impact on performance). This methodology enables accurate gate-level estimation of both area and power overheads introduced by HTs.